# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 2.3 On the Moreau Envelope Function Reformulation
Next, we show that $\nabla v_{\gamma}({x},{y})$ is continuous on $X\times\mathbb{R}^{m}$ by showing the continuity of $\theta^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. Let $\{(x^{k},y^{k})\}\subset X\times\mathbb{R}^{m}$ be any sequence such that $(x_{k},y_{k})\rightarrow(\bar{x},\bar{y})$. Since $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$, the sequence $\theta^{*}_{\gamma}(x_{k},y_{k})$ is bounded. Let $\bar{\theta}$ be an accumulation point of $\{\theta^{*}_{\gamma}(x_{k},y_{k})\}$. For any $\theta\in Y$, we have $f(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+g(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(x_{k},y_{k})-y_{k}\|^{2}\leq f(x_{k},\theta)+g(x_{k},\theta)+\frac{1}{2\gamma}\|\theta-y_{k}\|^{2}$. Taking the limit as $k\rightarrow\infty$, and using the continuity of $f$ and $g$, we obtain that for any $\theta\in Y$, $f(\bar{x},\bar{\theta})+g(\bar{x},\bar{\theta})+\frac{1}{2\gamma}\|\bar{\theta}-\bar{y}\|^{2}\leq f(\bar{x},\theta)+g(\bar{x},\theta)+\frac{1}{2\gamma}\|\theta-\bar{y}\|^{2}$. This implies $\bar{\theta}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, $\theta^{*}_{\gamma}(x_{k},y_{k})\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$, establishing the continuity of $\theta^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. ∎
@@
For any $y_{1},y_{2}\in\mathbb{R}^{m}$, let $\theta^{*}_{1}:=\theta_{\gamma}^{*}(x,y_{1})$ and $\theta^{*}_{2}:=\theta_{\gamma}^{*}(x,y_{2})$ be optimal solutions to problem $\min_{\theta\in Y}\,\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ with $y=y_{1}$ and $y_{2}$, respectively. From the first-order optimality conditions, we have
@@
$\leq$ $\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{2})/\gamma\right)\right)\big{\|}.$
Since $f(x,\theta)+\frac{1}{2\gamma}\|\theta-y_{1}\|^{2}$ is $1/\gamma$-strongly convex with respect to $\theta$ on $Y$, we obtain
@@
Let $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$, then
$\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{1})/\gamma\right)\right)\big{\|}^{2}$
$\leq$ $\left(1-s/\gamma\right)\|\theta^{*}_{1}-\theta^{*}_{2}\|^{2}.$
@@
$\|\theta^{*}_{1}-\theta^{*}_{2}\|\leq\sqrt{1-s/\gamma}\big{\|}\theta^{*}_{1}-\theta^{*}_{2}\big{\|}+s/\gamma\cdot\|y_{1}-y_{2}\|,$ (12)
@@
Next, following the approaches in [29, 27], we provide a characterization of solution quality for problem $({\rm VP})_{\gamma}^{\epsilon}$. We begin by introducing a constraint qualification condition that will be used in the analysis.
Definition 1.
Let $(\bar{x},\bar{y})\in X\times Y$. We say that the extended no nonzero abnormal multiplier constraint qualification (ENNAMCQ) holds at $(\bar{x},\bar{y})$ for problem $({\rm VP})_{\gamma}^{\epsilon}$ if either $f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})<\epsilon$ or $f(\bar{x},\bar{y})+g(\bar{x},\bar{y})-v_{\gamma}(\bar{x},\bar{y})\geq\epsilon$ but
$0\notin\nabla f(\bar{x},\bar{y})+\partial g(\bar{x},\bar{y})-\nabla v_{\gamma}(\bar{x},\bar{y})+\mathcal{N}_{X\times Y}(\bar{x},\bar{y}).$ (13)
For any $\epsilon>0$, as shown in Proposition 17 of [29], ENNAMCQ holds for problem $({\rm VP})_{\gamma}^{\epsilon}$ if the lower-level problem is strictly convex with respect to $y$.
Proposition 3.
Suppose that Assumptions 2 and 3 hold and $\gamma>0$. Then for any $\epsilon>0$ and $(\bar{x},\bar{y})\in X\times Y$, problem $({\rm VP})_{\gamma}^{\epsilon}$ satisfies ENNAMCQ at $(\bar{x},\bar{y})$, provided that the lower level problem is strictly convex for $\bar{x}$, i.e., $\varphi(\bar{x},y)+\delta_{Y}(y)$ is strictly convex with respect to $y$.
Having established the constraint qualification condition, we now define the concept of a stationary point for problem $({\rm VP})_{\gamma}^{\epsilon}$. This serves as a candidate for optimal solutions of the problem under the ENNAMCQ.
@@
Using 3, following [29, 27], we have the following necessary optimality condition.
Theorem 1.
Assume Assumptions 2 and 3 hold and $\gamma>0$. Let $(\bar{x},\bar{y})$ be a local optimal solution to problem ${\rm(VP)}_{\gamma}^{\epsilon}$ with $\epsilon>0$, and the lower level problem is strictly convex with respect to $y$ for $\bar{x}$. Then $(\bar{x},\bar{y})$ is a KKT point of problem ${\rm(VP)}_{\gamma}^{\epsilon}$.
@@ 3 Algorithm design
$\min_{x\in X,y\in Y}\psi_{p_{k}}(x,y):=\frac{1}{p_{k}}F(x,y)+f(x,y)+g(x,y)-v_{\gamma}(x,y),$
where $p_{k}$ is the penalty parameter at iteration $k$. Due the nonsmoothness of $g(x,y)$ with respect to $y$ and the availability of its proximal operator, it is natural to adopt a proximal alternating linearization method [52] to construct the algorithm. Given the current iterate $(x^{k},y^{k})$, linearizing the smooth components of $\psi_{p_{k}}(x,y)$ requires evaluating gradients, including $v_{\gamma}(x,y)$ as provided in 1. Calculating the exact gradient $\nabla v_{\gamma}(x^{k},y^{k})$ requires solving the proximal lower-level problem in (8) to obtain $\theta_{\gamma}^{*}(x^{k},y^{k})$. However, this solution typically does not have a closed-form expression, and solving it requires an iterative optimization solver, which can be computationally expensive. To address this, we propose an inexact gradient approximation by substituting $\theta_{\gamma}^{*}$ with an inexact estimate $\theta^{k}$.
The proposed algorithm alternates between updating $y$ and $x$ while refining the inexact approximation of $\theta_{\gamma}^{*}(x^{k},y^{k})$. Given a current iterate $(x^{k},y^{k})$ and $\theta^{k}$ for each $k=0,1,\ldots$, the next iterate $(x^{k+1},y^{k+1})$ and $\theta^{k+1}$ are updated as follows.
@@
$d_{y}^{k}:=\frac{1}{p_{k}}\nabla_{y}F\left(x^{k},y^{k}\right)+\nabla_{y}f\left(x^{k},y^{k}\right)-\frac{y^{k}-\theta^{k}}{\gamma},$ (14)
where $(y^{k}-\theta^{k})/{\gamma}$ is an inexact approximation of $\nabla_{y}v_{\gamma}(x^{k},y^{k})=({y^{k}}-\theta_{\gamma}^{*}(x^{k},y^{k}))/\gamma$ (see (9)). Using $d_{y}^{k}$, the variable $y$ is updated as
$y^{k+1}=\underset{y\in Y}{\mathrm{argmin}}~{}\left\langle d_{y}^{k},y-y^{k}\right\rangle+g(x^{k},y)+\frac{1}{2\beta_{k}}\left\|y-y^{k}\right\|^{2},$ (15)
@@
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k},\cdot)}\left(y^{k}-\beta_{k}d^{k}_{y}\right),$ (16)
@@
Find inexact approximation $\theta^{k+1/2}$. After updating $y^{k+1}$, we compute an inexact approximation $\theta^{k+1/2}$ for $\theta^{*}(x^{k},y^{k+1})$ by solving
@@
subject to the inexact criterion
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq s_{k},$ (18)
where $s_{k}$ controls the inexactness, and satisfies $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$. The prox-gradient residual $\mathcal{G}(\theta,x,y)$ quantifies the inexactness and is defined as
@@
where $\eta>0$. Note that $\mathcal{G}(\theta,x,y)=0$ if and only if $\theta=\theta^{*}(x,y)$.
@@
where $\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)-\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right)$ is an approximation to $\nabla_{x}v_{\gamma}(x^{k},y^{k})=\nabla_{x}f(x^{k},\theta^{*}(x^{k},y^{k}))-\nabla_{x}g(x^{k},\theta^{*}(x^{k},y^{k}))$ (see (9)). Using the direction $d_{x}^{k}$, we update $x$ with stepsize $\alpha_{k}>0$ as
@@
Find inexact approximation $\theta^{k+1}$. After updating $x^{k+1}$, we compute a new inexact approximation $\theta^{k+1}$ for $\theta^{*}(x^{k+1},y^{k+1})$ satisfying the inexact criterion
$\mathcal{G}(\theta^{k+1},x^{k+1},y^{k+1})\leq s_{k+1}.$ (21)
Update the penalty parameter $p_{k+1}$. To ensure the feasibility of the constraint $\varphi(x,y)-v_{\gamma}(x,y)\leq\epsilon$, we update the penalty parameter $p_{k}$ at each iteration. Evaluating the constraint violation requires computing $v_{\gamma}(x,y)$, which involves solving the proximal lower-level problem in (8). Instead, we use $\theta^{k+1}$ to update the penalty parameter. Specifically, we compute an approximation to the constraint violation as
@@
$t^{k+1}=\max\left\{\varphi(x^{k+1},y^{k+1})-\varphi(x^{k+1},\theta^{k+1})-\frac{1}{2\gamma}\|\theta^{k+1}-y^{k+1}\|^{2}-\epsilon,0)\right\}.$ (22)
The penalty parameter is then updated as
$p_{k+1}=\begin{cases}p_{k}+\varrho_{p},&\text{if}\quad\left\|(x^{k+1},y^{k+1})-(x^{k},y^{k})\right\|<c_{p}\min\left\{1/p_{k},{t_{k+1}}\right\},\\ p_{k},&\text{otherwise},\end{cases}$ (23)
where $\varrho_{p}$ and $c_{p}$ are positive constants controlling the penalty update.
@@
With these steps, we are now ready to present the proposed AGILS in 1.
@@
1:  Input: Initial iterates $x^{0},y^{0},\theta^{0}$, stepsizes $\alpha_{k},\beta_{k}$, relaxation parameter $\epsilon$, proximal parameter $\gamma$, penalty parameters $p_{0}$, $\varrho_{p}$, $c_{p}$, inexact parameters $\eta$, $s_{k}$, tolerance $tol$.
@@
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k,\cdot})}\left(y^{k}-\beta_{k}d^{k}_{y}\right).$

### v2

@@ 2.3 On the Moreau Envelope Function Reformulation
Next, we show that $\nabla v_{\gamma}({x},{y})$ is continuous on $X\times\mathbb{R}^{m}$ by showing the continuity of $\theta_{\gamma}^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. Let $\{(x_{k},y_{k})\}\subset X\times\mathbb{R}^{m}$ be any sequence such that $(x_{k},y_{k})\rightarrow(\bar{x},\bar{y})$. Since $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$, the sequence $\{\theta^{*}_{\gamma}(x_{k},y_{k})\}$ is bounded. Let $\bar{\theta}$ be an accumulation point of $\{\theta^{*}_{\gamma}(x_{k},y_{k})\}$. For any $\theta\in Y$, we have $f(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+g(x_{k},\theta^{*}_{\gamma}(x_{k},y_{k}))+\frac{1}{2\gamma}\|\theta^{*}_{\gamma}(x_{k},y_{k})-y_{k}\|^{2}\leq f(x_{k},\theta)+g(x_{k},\theta)+\frac{1}{2\gamma}\|\theta-y_{k}\|^{2}$. Taking the limit as $k\rightarrow\infty$, and using the continuity of $f$ and $g$, we obtain that for any $\theta\in Y$, $f(\bar{x},\bar{\theta})+g(\bar{x},\bar{\theta})+\frac{1}{2\gamma}\|\bar{\theta}-\bar{y}\|^{2}\leq f(\bar{x},\theta)+g(\bar{x},\theta)+\frac{1}{2\gamma}\|\theta-\bar{y}\|^{2}$. This implies $\bar{\theta}=\theta^{*}_{\gamma}(\bar{x},\bar{y})$. Therefore, $\theta^{*}_{\gamma}(x_{k},y_{k})\rightarrow\theta^{*}_{\gamma}(\bar{x},\bar{y})$, establishing the continuity of $\theta_{\gamma}^{*}(x,y)$ on $X\times\mathbb{R}^{m}$. ∎
@@
For any $y_{1},y_{2}\in\mathbb{R}^{m}$, let $\theta^{*}_{1}:=\theta_{\gamma}^{*}(x,y_{1})$ and $\theta^{*}_{2}:=\theta_{\gamma}^{*}(x,y_{2})$ be optimal solutions to the problem $\min_{\theta\in Y}\,\varphi(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ with $y=y_{1}$ and $y_{2}$, respectively. From the first-order optimality conditions, we have
@@
$\leq$ $\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{2})/\gamma\right)\right)\big{\|}$
$\leq$ $\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{1})/\gamma\right)\right)\big{\|}$
$+\frac{s}{\gamma}\|y_{1}-y_{2}\|,$
where the last inequality follows from the decomposition $s(\theta^{*}_{2}-y_{2})/\gamma=s(\theta^{*}_{2}-y_{1})/\gamma+s(y_{1}-y_{2})/\gamma$ and the triangle inequality.
Since $\nabla_{y}f(x,\cdot)$ is $L_{f_{y}}$-Lipschitz continuous on $Y$, it holds that
$\left\|\nabla_{y}f(x,\theta_{1}^{*})+(\theta_{1}^{*}-y_{1})/\gamma-\nabla_{y}f(x,\theta_{2}^{*})-(\theta_{2}^{*}-y_{1})/\gamma\right\|\leq\left(L_{f_{y}}+1/\gamma\right)\|\theta_{1}^{*}-\theta_{2}^{*}\|.$
Furthermore, since $f(x,\theta)+\frac{1}{2\gamma}\|\theta-y_{1}\|^{2}$ is $1/\gamma$-strongly convex with respect to $\theta$ on $Y$, we obtain
@@
Now, let us square the first term in the last line of (11). Applying the above bounds and choosing $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$, we obtain
$\quad\,\left\|\left(\theta_{1}^{*}-s\left(\nabla_{y}f(x,\theta_{1}^{*})+(\theta_{1}^{*}-y_{1})/\gamma\right)\right)-\left(\theta_{2}^{*}-s\left(\nabla_{y}f(x,\theta_{2}^{*})+(\theta_{2}^{*}-y_{1})/\gamma\right)\right)\right\|^{2}$
$=\|\theta_{1}^{*}-\theta_{2}^{*}\|^{2}+s^{2}\left\|\nabla_{y}f(x,\theta_{1}^{*})+(\theta_{1}^{*}-y_{1})/\gamma-\nabla_{y}f(x,\theta_{2}^{*})-(\theta_{2}^{*}-y_{1})/\gamma\right\|^{2}$
$\quad-2s\left\langle\nabla_{y}f(x,\theta_{1}^{*})+(\theta_{1}^{*}-y_{1})/\gamma-\nabla_{y}f(x,\theta_{2}^{*})-(\theta_{2}^{*}-y_{1})/\gamma,\ \theta_{1}^{*}-\theta_{2}^{*}\right\rangle$
$\leq\left(1+s^{2}\left(L_{f_{y}}+1/\gamma\right)^{2}-2s/\gamma\right)\|\theta_{1}^{*}-\theta_{2}^{*}\|^{2}$
$=\left(1-s/\gamma\right)\|\theta_{1}^{*}-\theta_{2}^{*}\|^{2}.$
Consequently, we have
$\big{\|}\left(\theta^{*}_{1}-s\left(\nabla_{y}f(x,\theta^{*}_{1})+(\theta^{*}_{1}-y_{1})/\gamma\right)\right)-\left(\theta^{*}_{2}-s\left(\nabla_{y}f(x,\theta^{*}_{2})+(\theta^{*}_{2}-y_{1})/\gamma\right)\right)\big{\|}$
$\leq$ $\sqrt{1-s/\gamma}\big{\|}\theta^{*}_{1}-\theta^{*}_{2}\big{\|}.$
@@
$\|\theta^{*}_{1}-\theta^{*}_{2}\|\leq\sqrt{1-s/\gamma}\big{\|}\theta^{*}_{1}-\theta^{*}_{2}\big{\|}+s/\gamma\cdot\|y_{1}-y_{2}\|,$
@@
This section concludes with the concept of a stationary point for problem $({\rm VP})_{\gamma}^{\epsilon}$. This stationary point, under certain constraint qualifications detailed in [5, 31], can be considered a candidate for an optimal solution.
@@ 3 Algorithm design
$\min_{x\in X,y\in Y}\psi_{p_{k}}(x,y):=\frac{1}{p_{k}}F(x,y)+f(x,y)+g(x,y)-v_{\gamma}(x,y),$ (12)
where $p_{k}$ is the penalty parameter at iteration $k$. Due to the nonsmoothness of $g(x,y)$ with respect to $y$ and the availability of its proximal operator, it is natural to adopt a proximal alternating linearization method [14] to construct the algorithm. Given the current iterate $(x^{k},y^{k})$, linearizing the smooth components of $\psi_{p_{k}}(x,y)$ requires evaluating gradients, including $\nabla v_{\gamma}(x,y)$ as provided in Proposition 1. Calculating the exact gradient $\nabla v_{\gamma}(x^{k},y^{k})$ requires solving the proximal lower-level problem in (8) to obtain $\theta_{\gamma}^{*}(x^{k},y^{k})$. However, this solution typically does not have a closed-form expression, and solving it requires an iterative optimization solver, which can be computationally expensive. To address this, we propose an inexact gradient approximation by substituting $\theta_{\gamma}^{*}$ with an inexact estimate $\theta^{k}$.
The proposed algorithm alternates between updating $y$ and $x$ while refining the inexact approximation of $\theta_{\gamma}^{*}(x^{k},y^{k})$. At each iteration $k=0,1,\ldots$, given the current iterate $(x^{k},y^{k})$, along with the feasibility-corrected iterate $\tilde{y}^{k}$ and corresponding $\tilde{\theta}^{k}$(constructed from $y^{k}$ and $\theta^{k}$ via the procedure detailed below), the next iterate $(x^{k+1},y^{k+1})$ and the refined approximation $\theta^{k+1}$ are updated as follows.
@@
$d_{y}^{k}:=\frac{1}{p_{k}}\nabla_{y}F\left(x^{k},\tilde{y}^{k}\right)+\nabla_{y}f\left(x^{k},\tilde{y}^{k}\right)-\frac{\tilde{y}^{k}-\tilde{\theta}^{k}}{\gamma},$ (13)
where $(\tilde{y}^{k}-\tilde{\theta}^{k})/{\gamma}$ is an inexact approximation of $\nabla_{y}v_{\gamma}(x^{k},\tilde{y}^{k})=({\tilde{y}^{k}}-\theta_{\gamma}^{*}(x^{k},\tilde{y}^{k}))/\gamma$ (see (9)). Using $d_{y}^{k}$, the variable $y$ is updated as
$y^{k+1}=\underset{y\in Y}{\mathrm{argmin}}~{}\left\langle d_{y}^{k},y-\tilde{y}^{k}\right\rangle+g(x^{k},y)+\frac{1}{2\beta_{k}}\left\|y-\tilde{y}^{k}\right\|^{2},$ (14)
@@
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k},\cdot)}\left(\tilde{y}^{k}-\beta_{k}d^{k}_{y}\right),$ (15)
@@
Find inexact approximation $\theta^{k+1/2}$. After updating $y^{k+1}$, we compute an inexact approximation $\theta^{k+1/2}$ for $\theta_{\gamma}^{*}(x^{k},y^{k+1})$ by solving
@@
subject to either an absolute or a relative inexactness criterion
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq s_{k},\quad\text{or}\quad\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq\tau_{k}\mathcal{G}(\theta^{k-1},x^{k-1},y^{k-1}).$ (17)
The parameters $s_{k}$ and $\tau_{k}$ control the solution’s inexactness, and are required to satisfy $\sum_{k=0}^{\infty}s_{k}^{2}<\infty$ and $\tau_{k}:=\tau_{0}/(k+1)^{p_{\tau}}$ with constants $\tau_{0}>0$ and $p_{\tau}>0$. The prox-gradient residual $\mathcal{G}(\theta,x,y)$ quantifies the inexactness and is defined as
@@
where $\eta>0$. Note that $\mathcal{G}(\theta,x,y)=0$ if and only if $\theta=\theta_{\gamma}^{*}(x,y)$.
Remark 2.
A similar $\theta$-update based on solving the proximal subproblem (16) is also employed by MEHA in [42]. However, MEHA performs only a single proximal gradient step to update $\theta$. In contrast, our method uses an inner loop governed by the inexactness criteria (17). As we will show, this strategy yields the advantage of a larger and explicitly defined range for the step sizes.
@@
where $\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)+\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right)$ is an approximation to $\nabla_{x}v_{\gamma}(x^{k},y^{k})=\nabla_{x}f(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k}))+\nabla_{x}g(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k}))$ (see (9)). Using the direction $d_{x}^{k}$, we update $x$ with stepsize $\alpha_{k}>0$ as
@@
Find inexact approximation $\theta^{k+1}$. After updating $x^{k+1}$, a new inexact approximation $\theta^{k+1}$ of $\theta^{*}_{\gamma}(x^{k+1},y^{k+1})$ is computed by approximately solving
@@
$\theta^{k+1}\approx\underset{y\in Y}{\mathrm{argmin}}\left\{f(x^{k+1},\theta)+g(x^{k+1},\theta)+\frac{1}{2\gamma}\left\|\theta-y^{k+1}\right\|^{2}\right\},$
such that $\theta^{k+1}$ satisfies either an absolute or a relative inexactness criterion
$\mathcal{G}(\theta^{k+1},x^{k+1},y^{k+1})\leq s_{k+1},\quad\text{or}\quad\mathcal{G}(\theta^{k+1},x^{k+1},y^{k+1})\leq\tau_{k+1}\mathcal{G}(\theta^{k},x^{k},y^{k}).$ (20)
Remark 3.
The parameters $s_{k+1}$ and $\tau_{k+1}$ used in the inexact criteria (20) for determining $\theta^{k+1}$ can be chosen from sequences independent of those used for $\theta^{k+1/2}$, provided they satisfy the previously stated requirements.
Penalty parameter update and feasibility correction. To encourage feasibility of the constraint $\varphi(x,y)-v_{\gamma}(x,y)\leq\epsilon$, the penalty parameter $p_{k}$ is updated at each iteration. Evaluating the constraint violation requires computing $v_{\gamma}(x,y)$, which involves solving the proximal lower-level problem in (8). Instead, we use $\theta^{k+1}$ to estimate the constraint violation as
$t^{k+1}=\max\left\{\varphi(x^{k+1},y^{k+1})-\varphi(x^{k+1},\theta^{k+1})-\frac{1}{2\gamma}\|\theta^{k+1}-y^{k+1}\|^{2}-\epsilon,0\right\}.$ (21)
The penalty parameter $p_{k}$ is increased to $p_{k+1}=p_{k}+\varrho_{p}$ (for some $\varrho_{p}>0$) only when the change in the iterates, $\Delta_{k+1}:=\left\|(x^{k+1},y^{k+1})-(x^{k},\tilde{y}^{k})\right\|$, is small relative to the penalty parameter (i.e., $\Delta_{k+1}<c_{p}/p_{k}$ for some $c_{p}>0$), while the estimated constraint violation $t^{k+1}$ is large relative to this change (i.e., $t^{k+1}>\Delta_{k+1}/c_{p}$). These two conditions can be compactly written as
$\Delta_{k+1}<c_{p}\min\left\{1/p_{k},{t^{k+1}}\right\}.$ (22)
However, simply increasing $p_{k}$ does not always guarantee feasibility. It can sometimes cause the algorithm to stagnate near an infeasible stationary point of the constraint function $\varphi(x,y)-v_{\gamma}(x,y)$. To mitigate this, we use the bilevel optimization structure, that the feasible region is the solution set of the convex lower-level problem (2). We introduce a feasibility correction procedure designed to help the iterates escape such stationary points.
The feasibility correction procedure aims to generate a corrected iterate $\tilde{y}^{k+1}$ from $y^{k+1}$. It is based on the observation that if $(x^{k+1},y^{k+1})$ is feasible, then $y^{k+1}$ should be a reasonably accurate solution to the lower-level problem (2) for $x=x^{k+1}$, implying that $\|\nabla_{y}v_{\gamma}(x^{k+1},y^{k+1})\|$ should be small. We use the magnitude of $(y^{k+1}-\theta^{k+1})/\gamma$ to approximate $\|\nabla_{y}v_{\gamma}(x^{k+1},y^{k+1})\|$. The feasibility correction is triggered if:
$\left\|y^{k+1}-\theta^{k+1}\right\|\geq\frac{c_{y}\gamma}{p_{k}},$ (23)
where $c_{y}>0$. Condition (23) indicates that the iterate $(x^{k+1},y^{k+1})$ might be approaching an undesirable stationary point of the constraint function.
If condition (23) holds, we attempt to correct $y^{k+1}$ by finding a candidate iterate $\tilde{y}^{k+1}$ as an inexact solution to the lower-level problem (2) with $x=x^{k+1}$ satisfying
$\left\|\tilde{y}^{k+1}-\operatorname{Prox}_{\tilde{g}(x^{k+1},\cdot)}\left(\tilde{y}^{k+1}-\nabla_{y}f(x^{k+1},\tilde{y}^{k+1})\right)\right\|\leq\frac{c_{\tilde{y}}}{p_{k}}\Delta_{k+1},$ (24)
where $c_{\tilde{y}}>0$. However, unconditionally accepting such a $\tilde{y}^{k+1}$ could adversely affect the algorithm’s convergence. Therefore, to preserve convergence, $\tilde{y}^{k+1}$ is accepted only if it satisfies the following descent condition
$\frac{1}{p_{k}}F(x^{k+1},\tilde{y}^{k+1})+\varphi(x^{k+1},\tilde{y}^{k+1})-\left(\varphi(x^{k+1},\tilde{\theta}^{k+1})+\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}\right)$ (25)
$\leq$ $\frac{1}{p_{k}}F(x^{k+1},y^{k+1})+\varphi(x^{k+1},y^{k+1})-\left(\varphi(x^{k+1},\theta^{k+1})+\frac{1}{2\gamma}\|\theta^{k+1}-y^{k+1}\|^{2}\right),$
where $\tilde{\theta}^{k+1}$ is an approximation to $\theta^{*}_{\gamma}(x^{k+1},\tilde{y}^{k+1})$ satisfying the inexactness criterion
$\mathcal{G}(\tilde{\theta}^{k+1},x^{k+1},\tilde{y}^{k+1})\leq s_{k+1},\quad\text{or}\quad\mathcal{G}(\tilde{\theta}^{k+1},x^{k+1},\tilde{y}^{k+1})\leq\tau_{k+1}\mathcal{G}(\theta^{k},x^{k},y^{k}).$ (26)
If the descent condition (25) holds, the algorithm adopts the candidate $\tilde{y}^{k+1}$ along with its corresponding $\tilde{\theta}^{k+1}$, and keep the penalty parameter unchanged, i.e., $p_{k+1}=p_{k}$. Otherwise, the correction is discarded, meaning $\tilde{y}^{k+1}=y^{k+1}$ and $\tilde{\theta}^{k+1}=\theta^{k+1}$ and the penalty parameter is increased, i.e., $p_{k+1}=p_{k}+\varrho_{p}$. The algorithm then proceeds with the selected pair $(\tilde{y}^{k+1},\tilde{\theta}^{k+1})$. A detailed description of this feasibility correction procedure is presented in Algorithm 2.
@@
With these components established, we are now ready to present the proposed AGILS algorithm in Algorithm 1.
@@
1:  Input: Initial iterates $x^{0},y^{0},\tilde{\theta}^{0},\tilde{y}^{0}$, set $(x^{-1},y^{-1},\tilde{\theta}^{-1})=(x^{0},y^{0},\tilde{\theta}^{0})$, stepsizes $\alpha_{k},\beta_{k}$, relaxation parameter $\epsilon$, proximal parameter $\gamma$, penalty parameters $p_{0}$, $\varrho_{p}$, $c_{p}$, inexact parameters $\eta$, $s_{k}$, $\tau_{k}$, tolerance $tol$.
@@
$y^{k+1}=\operatorname{Prox}_{\beta_{k}\tilde{g}(x^{k,\cdot})}\left(\tilde{y}^{k}-\beta_{k}d^{k}_{y}\right).$